# TASAR: Transfer-based Attack on Skeletal Action Recognition

### v1

@@ Abstract
Skeletal sequences, as well-structured representations of human behaviors, are crucial in Human Activity Recognition (HAR). The transferability of adversarial skeletal sequences enables attacks in real-world HAR scenarios, such as autonomous driving, intelligent surveillance, and human-computer interactions. However, existing Skeleton-based HAR (S-HAR) attacks exhibit weak adversarial transferability and, therefore, cannot be considered true transfer-based S-HAR attacks. More importantly, the reason for this failure remains unclear. In this paper, we study this phenomenon through the lens of loss surface, and find that its sharpness contributes to the poor transferability in S-HAR. Inspired by this observation, we assume and empirically validate that smoothening the rugged loss landscape could potentially improve adversarial transferability in S-HAR. To this end, we propose the first Transfer-based Attack on Skeletal Action Recognition, TASAR. TASAR explores the smoothed model posterior without re-training the pre-trained surrogates, which is achieved by a new post-train Dual Bayesian optimization strategy. Furthermore, unlike previous transfer-based attacks that treat each frame independently and overlook temporal coherence within sequences, TASAR incorporates motion dynamics into the Bayesian attack gradient, effectively disrupting the spatial-temporal coherence of S-HARs. To exhaustively evaluate the effectiveness of existing methods and our method, we build the first large-scale robust S-HAR benchmark, comprising 7 S-HAR models, 10 attack methods, 3 S-HAR datasets and 2 defense models. Extensive results demonstrate the superiority of TASAR. Our benchmark enables easy comparisons for future studies, with the code available in the supplementary material.
Introduction
Human Activity Recognition (HAR) has a wide range of application scenarios, such as human-computer interactions, bio-mechanics, and virtual reality (Sun et al. 2022; Guo et al. 2024a). Skeleton sequences has been widely used in HAR (Yan, Xiong, and Lin 2018), because skeleton data is a well-structured data representation of human behaviors and robust to lighting, occlusion, and view angles. Recent research (Wang et al. 2021; Diao et al. 2021) demonstrates that S-HARs are less robust than expected when adversarial perturbations are added to the testing data, causing skeletal classifiers to produce incorrect outputs. Previous skeleton-based attacks are mainly developed under the white-box setting (Liu, Akhtar, and Mian 2020; Tanaka, Kera, and Kawamoto 2022), where the attacker knows the structure and parameters of a given model, or under the query-based setting (Diao et al. 2021; Kang et al. 2023a), where the adversary can request numerous queries against the target model (Diao et al. 2024a). But both settings are impractical in real-world HAR scenarios, such as autonomous driving (Guo et al. 2024b), intelligent surveillance (Garcia-Cobo and SanMiguel 2023), and human-computer interactions (Wang et al. 2020), in which the white-box information and a large number of queries are not attainable.
In contrast, transfer-based attacks, which craft adversarial examples from surrogate models and then transfer them to target black-box models, present a more realistic threat under the free-query black-box setting. Although adversarial transferability has been widely studied across various tasks (Huang et al. 2023; Dong et al. 2018; Diao et al. 2024b), research on S-HAR remains limited. Recent studies have attempted to apply white-box S-HAR attacks against black-box models via surrogate models (Wang et al. 2021; Liu, Akhtar, and Mian 2020). However, their transferability is low and highly sensitive to the surrogate choice (Wang et al. 2023; Lu et al. 2023). Consequently, existing S-HAR attacks cannot be regarded as transferable. Similarly, previous transfer-based attacks (Dong et al. 2018; Xiong et al. 2022) are successful on image data, but show poor transferability when applied to skeletal motion. Contrary to the common belief that adversarial examples transfer well across different model architectures and parameters (Liu et al. 2016), existing attacks do not exhibit transferability on S-HAR, raising doubt on the usefulness of adversarial transferability in this domain (Lu et al. 2023). More importantly, the reason for this failure remains unclear.
To study this phenomenon, we begin by examining the factors that hinder adversarial transferability in S-HARs. Our preliminary experiments reveal that adversarial transferability is highly sensitive to the chosen surrogates (Tab. 1). This finding motivates us to further explore the differences between surrogate models from the view of loss surface smoothness, as previous research has demonstrated that the smoothness of the loss surface significantly impacts adversarial transferability (Wu and Zhu 2020; Qin et al. 2022). As shown in Fig. 2, surrogates with a smoothed loss landscape have higher transferability than surrogates with a sharp loss landscape. Therefore, unlike most existing attacks that focus on the backpropagation process, we argue that the S-HAR attack should prioritize smoothening the surrogate’s landscape in the training process.
Figure 1: A high-level illustration of our proposed method.
@@ Introduction
Bayesian neural networks have been shown to exhibit lower sharpness and better generalization ability (Blundell et al. 2015; Maddox et al. 2019). Therefore, attacking the entire model distribution is more likely to generalize for S-HARs (Li et al. 2023; Gubri et al. 2022). However, it is not straightforward to design such a transferable Bayesian attack for S-HAR. First, skeletal classifiers contain at least several millions of parameters. Sampling from the posterior distribution for such complex models can be computationally expensive and consume a heavy memory footprint. Second, most prior transferable attacks are specifically designed for static data, e.g. images. However, most S-HAR models learn the spatial-temporal features because skeletal data contains rich motion dynamics. Naive adaptation of the Bayesian attack may ignore the spatial-temporal coherence when issuing attacks. How to consider the motion dynamics in Bayesian attacks has not been explored.
To tackle these challenges, we propose the first Transfer-based Attack specifically designed for Skeletal Action Recognition, TASAR, based on a Post-train Dual Bayesian optimization that significantly smoothens the rugged landscape (see Fig. 2). Our post-train Bayesian strategy freezes the pre-trained surrogate and appends small extra Bayesian components behind it, to turn a single surrogate into a Bayesian one without the need for re-training, thereby speeding up the training process and avoiding a heavy memory footprint. Next, we integrate the temporal motion gradient in a Bayesian manner to disrupt the temporal coherence between sequences over time. A high-level illustration of our method is presented in Fig. 1. Our contributions are summarized as follows.
  • We systematically investigate the previously unknown reason for low adversarial transferability in S-HAR through the lens of loss surface smoothness, and introduce the first transfer-based attack on S-HAR to address this issue.
  • We propose a novel post-train Dual Bayesian Motion attack, that explores the smoothed model posterior and considers the spatial-temporal coherence from a Bayesian perspective, without requiring re-training of the pre-trained surrogate model.
  • We build the first comprehensive robust S-HAR evaluation benchmark RobustBenchHAR. RobustBenchHAR consists of 7 S-HAR models with diverse GCN structures and more recent Transformer structures, 10 attack methods, 3 datasets and 2 defense methods. Extensive experiments on this benchmark demonstrate the superiority and generalizability of TASAR.
Related Work
@@
Skeleton-Based Human Action Recognition. Early S-HAR research employed convolutional neural networks (CNNs) (Ali et al. 2023) and recurrent neural networks (RNNs) (Du, Wang, and Wang 2015) to extract motion features in the spatial domain and temporal domains, respectively. However, skeleton data, inherently a topological graph, poses challenges for feature representation using traditional CNNs and RNNs. Recent advances with graph convolutional networks (GCNs) (Kipf and Welling 2016) have improved performance by modeling skeletons as topological graphs, with nodes corresponding to joints and edges to bones (Yan, Xiong, and Lin 2018). Subsequent improvements in graph designs and network architectures include two-stream adaptive GCN (2s-AGCN) (Shi et al. 2019b), directed acyclic GCN (DGNN) (Shi et al. 2019a), multi-scale GCN (MS-G3D) (Liu et al. 2020), channel-wise topology refinement (CTR-GCN) (Chen et al. 2021) and auxiliary feature refinement (FR-HEAD) (Zhou, Liu, and Wang 2023). Alongside advancements in GCN-based models, recent studies are exploring temporal Transformer structures for S-HARs (Do and Kim 2024; Qiu et al. 2022; Guo et al. 2024b), but their vulnerability remains unexplored. This work is the first to assess the adversarial robustness of Transformer-based S-HARs.
Adversarial Attacks on S-HAR. Adversarial attacks (Szegedy et al. 2013) highlight the susceptibility of deep neural networks and have been applied across different data types. Recently, attacks on S-HAR have garnered increasing attention. CIASA (Liu, Akhtar, and Mian 2020) proposes a constrained iterative attack via GAN (Goodfellow et al. 2014) to regularize the adversarial skeletons. SMART (Wang et al. 2021) proposes a perception loss gradient. Tanaka et al. (Tanaka, Kera, and Kawamoto 2022) suggest only perturbing skeletal lengths, and evaluate robustness via Fourier analysis (Tanaka, Kera, and Kawamoto 2024). These methods are white-box attacks, requiring full knowledge of the victim model. In contrast, BASAR (Diao et al. 2021, 2024a) proposes motion manifold searching to achieve the query-based black-box attack. FGDA-GS (Kang et al. 2023b) estimates gradient signs to further reduce query numbers.
@@
Compared to white-box and query-based attacks, transfer-based attacks (Liu et al. 2016) pose a more practical threat as real-world HAR scenarios typically cannot access white-box information or extensive querying. While existing white-box S-HAR attacks (Wang et al. 2021; Liu, Akhtar, and Mian 2020) can be adapted for transfer-based scenarios, they suffer from low transferability and sensitivity to surrogate choices. Lu et al. (Lu et al. 2023) proposes a no-box attack for S-HAR but also lacks transferability. Numerous transfer-based attacks, including gradient-based (Lin et al. 2019), input transformation (Xie et al. 2019), and ensemble-based methods (Xiong et al. 2022; Li et al. 2023), exhibit high transferability across various tasks but struggle in skeletal data (Lu et al. 2023). Therefore, there is an urgent need to develop a transferable attack for skeleton-based action recognition.
Methodology
Preliminaries
@@
We denote a clean motion $\mathbf{x}\in\mathcal{X}$ and its corresponding label $y\in\mathcal{Y}$. Given a surrogate action recognizer $f_{\theta}$ parametrized by $\theta$, $f_{\theta}$ is trained to map a motion $\mathbf{x}$ to a predictive distribution $p(y\mid\mathbf{x},\theta)$. The white-box attack can be optimized by minimizing the predictive probability:
$\mathop{\arg\min}\limits_{\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon}p(y\mid\tilde{\mathbf{x}},\theta)$ (1)
@@
where $\tilde{\mathbf{x}}$ is the adversarial example and $\epsilon$ is the perturbation budget. $\|\cdot\|_{p}$ is the $l_{p}$ norm distance. The procedure of transfer-based attack is firstly crafting the adversarial example $\tilde{\mathbf{x}}$ by attacking the surrogate model, then transferring $\tilde{\mathbf{x}}$ to attack the unseen target model. In Eq. 1, since the transferable adversarial examples are optimized against one surrogate model, the adversarial transferability heavily relies on the surrogate model learning a classification boundary similar to that of the unknown target model. While possible for image classification, it proves unrealistic for S-HAR (Wang et al. 2023; Lu et al. 2023).
Motivation
Figure 2: Comparison of loss landscapes of trained models. This visualization method was proposed by (Li et al. 2018), which approximates neural network loss functions in low-dimensional subspaces of the model’s parameter space. The $x$ and $y$ axis represent two random direction vectors. The $z$ axis represents the loss. PB means the post-train Bayesian optimization, and P-DB means the improved post-train Dual Bayesian optimization. The loss landscape optimized by post-train Dual Bayesian is significantly smoother than those of vanilla post-train Bayesian and standard training. More visualizations can be found in Appendix B.
@@
Existing S-HAR attacks have shown outstanding white-box attack performance but exhibit low transferability (Wang et al. 2023). Similarly, previous transfer-based attacks (Dong et al. 2018; Xiong et al. 2022), successful on image data, also show poor transferability when applied to skeletal motion (Lu et al. 2023). Naturally, two questions occur to us: (1) Why do existing adversarial attacks fail to exhibit transferability in skeletal data? (2) Do transferable adversarial examples truly exist in S-HAR?
@@
To answer these questions, we start by generating adversarial examples using various surrogate skeletal recognizers and then evaluate their adversarial transferability. Obviously, in Tab. 1, the transferability is highly sensitive to the chosen surrogates, e.g. CTR-GCN (Chen et al. 2021) as the surrogate exhibits higher transferability than ST-GCN (Yan, Xiong, and Lin 2018). This observation motivates us to further investigate the differences between surrogate models. Previous research (Wu and Zhu 2020; Qin et al. 2022) has proven that adversarial examples generated by surrogate models with a less smooth loss landscape are unlikely to transfer across models. Therefore, we investigate the smoothness of the loss landscape across different surrogate models. In Fig. 2, We visualize the loss landscape of ST-GCN and CTR-GCN trained on the skeletal dataset NTU-60 (Shahroudy et al. 2016) and compare their smoothness to the ResNet-18 (He et al. 2016) trained on CIFAR-10 (Krizhevsky, Hinton et al. 2009). More landscape visualizations can be found in Appendix B. By analyzing the loss surface smoothness, we have two findings: (1) The loss surface of models trained on skeletal data is much sharper than those trained on image data, leading to a relatively low transferability. This suggests that adversarial examples within a sharp local region are less likely to transfer across models in S-HAR, potentially explaining our first question. (2) CTR-GCN has a flatter loss landscape compared to ST-GCN, making it a more effective surrogate for higher transferability. Consequently, we argue that using a surrogate with a smoothed loss landscape will significantly enhance adversarial transferability in S-HAR.
@@
In this work, motivated by evidence that Bayesian neural networks (BNNs) exhibit low sharpness and good generalization (Blundell et al. 2015; Maddox et al. 2019), we aim to construct a Bayesian surrogate by sampling from the model posterior space to smoothen the rugged loss landscape. From a Bayesian perspective, Eq. 1 can be reformulated by approximately minimizing the Bayesian posterior predictive distribution:
@@
$\underset{\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon}{\arg\min}\,p(y\mid\tilde{\mathbf{x}},\mathcal{D})=\underset{\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon}{\arg\min}\,\mathbb{E}_{\theta\sim p(\theta\mid\mathcal{D})}p\left(y\mid\tilde{\mathbf{x}},\theta\right)$ (2)
@@
A Post-train Bayesian Perspective on Attack
Unfortunately, directly sampling from the posterior distribution of skeletal classifiers is not a straightforward task due to several factors. First, directly sampling the posterior is intractable for large-scale skeletal classifiers. Although approximate methods such as MCMC sampling (Welling and Teh 2011) or variational inference (Blei, Kucukelbir, and McAuliffe 2017) are possible, sampling is prohibitively slow and resource-intensive due to the high dimensionality of the sampling space, which typically involves at least several million parameters in skeletal classifiers. In addition, skeletal classifiers normally contain a large number of parameters and are pre-trained on large-scale datasets (Liu et al. 2019). Consequently, it is not practical for end-users to re-train the surrogate in a Bayesian manner, as the training process is time-consuming.

### v5

@@ Abstract
Skeletal sequence data, as a widely employed representation of human actions, are crucial in Human Activity Recognition (HAR). Recently, adversarial attacks have been proposed in this area, which exposes potential security concerns, and more importantly provides a good tool for model robustness test. Within this research, transfer-based attack is an important tool as it mimics the real-world scenario where an attacker has no knowledge of the target model, but is under-explored in Skeleton-based HAR (S-HAR). Consequently, existing S-HAR attacks exhibit weak adversarial transferability and the reason remains largely unknown. In this paper, we investigate this phenomenon via the characterization of the loss function. We find that one prominent indicator of poor transferability is the low smoothness of the loss function. Led by this observation, we improve the transferability by properly smoothening the loss when computing the adversarial examples. This leads to the first Transfer-based Attack on Skeletal Action Recognition, TASAR. TASAR explores the smoothened model posterior of pre-trained surrogates, which is achieved by a new post-train Dual Bayesian optimization strategy. Furthermore, unlike existing transfer-based methods which overlook the temporal coherence within sequences, TASAR incorporates motion dynamics into the Bayesian attack, effectively disrupting the spatial-temporal coherence of S-HARs. For exhaustive evaluation, we build the first large-scale robust S-HAR benchmark, comprising 7 S-HAR models, 10 attack methods, 3 S-HAR datasets and 2 defense models. Extensive results demonstrate the superiority of TASAR. Our benchmark enables easy comparisons for future studies, with the code available in the https://github.com/yunfengdiao/Skeleton-Robustness-Benchmark.
[…]
1 Introduction
S-HAR has been an important research topic in computer vision. Recently, S-HAR classifiers have been found to be susceptible to adversarial attack (Wang et al., 2021; Diao et al., 2021), suggesting adversarial attack potentially provides a useful tool for robustness tests for S-HAR classifiers. But not all attacks are equally practical. Existing S-HAR attacks are mainly proposed under white-box settings (Liu et al., 2020a; Tanaka et al., 2022), where the attacker has full access to the victim model’s architecture, weights, and training details, or under query-based black-box settings (Diao et al., 2021; Kang et al., 2023b), where the attacker can make numerous queries (Diao et al., 2024a). However, neither approaches are impractical in real-world scenarios (e.g. autonomous driving (Guo et al., 2024), intelligent surveillance (Garcia-Cobo & SanMiguel, 2023) and human-computer interactions (Wang et al., 2020)), where either accessing the victim model or numerous queries is not attainable. Therefore, transfer-based attack, i.e. generating adversarial examples by attacking a surrogate model and then transfer them to target black-box models, is proposed as a promising alternative (Dong et al., 2018; Wang et al., 2021).
However, current transfer-based attack on S-HAR is far from ideal due to their generally poor and unreliable performance. Recently, few studies have attempted to apply white-box S-HAR attacks against black-box models via surrogate models (Wang et al., 2021; Liu et al., 2020a). However, results show that their transfer success rate is highly determined by the specific choice of the S-HAR surrogate, so that its general adversarial transferability is low (Wang et al., 2023; Lu et al., 2023), also referred to as low/weak transferability. Although similar research in other fields (Dong et al., 2018; Huang et al., 2023; Diao et al., 2024b) has achieved success, a direct application of them on S-HAR still shows low transferability, raising doubt on the usefulness of adversarial transferability in this domain (Lu et al., 2023). More importantly, the reason for this failure remains unclear.
We begin by investigating the underlying causes of the low transferability in S-HAR attacks. By first systematically investigating the sensitivity of attack transferability on the choice of surrogates, we compare the loss surface smoothness of the surrogates, inspired by (Wu & Zhu, 2020; Qin et al., 2022). A visual comparison is shown in Figure 2, which gives a clear indication of high correlations between loss smoothness and transferability. Consequently, we argue that the transfer-based S-HAR attack should smoothen the surrogate’s loss during training. Various strategies aim to achieve smoother loss landscapes, via e.g. regularization (Zhao et al., 2022; Foret et al., 2021) or Bayesian learning (Izmailov et al., 2018; Nguyen et al., 2024; Maddox et al., 2019). We explore the latter and use Bayesian Neural Networks (BNNs). This is because BNNs tend to have smooth loss landscapes (Blundell et al., 2015; Izmailov et al., 2018; Nguyen et al., 2024). More importantly, it enables us to attack the whole distributions of models, i.e. Bayesian attacks, which has been proven to enhance the transferability in other fields (Li et al., 2023; Gubri et al., 2022).
@@ 1 Introduction
However, it is not straightforward to design such a transferable Bayesian attack for S-HAR. First, attacking a distribution of models requires sampling from the posterior distribution. But S-HAR classifiers contain at least several millions of parameters (Liu et al., 2020b), which makes sampling computationally expensive. Second, most prior transferable attacks are specifically designed for static data, e.g. images. However, most S-HAR models learn the spatial-temporal features because skeletal data contains rich motion dynamics. A naive adaptation of them ignores the spatial-temporal coherence during attack, leading to either lower transferability or excessive attack which raises suspicion. How to incorporate the motion dynamics in Bayesian attacks has not been explored.
Figure 1: A high-level illustration of our proposed method. Results marked with a ‘check mark’ ($\surd$) indicate superior performance compared to those marked with a ‘cross’ ($\times$). Spatial attack: treats each frame independently. Spatial-temporal Attack: integrates temporal motion gradients to disrupt the spatial-temporal coherence of S-HAR models.
To tackle these challenges, we propose the first Transfer-based Attack specifically designed for Skeletal Action Recognition, TASAR, with key novelties shown in Figure 1. First, our post-train Bayesian strategy keeps a pre-trained surrogate intact by appending lightweight Bayesian components behind it, without the need for re-training of the pre-trained surrogate. Second, we propose a novel dual Bayesian optimization for smoothed posterior sampling, which effectively smoothens the rugged loss surface. Finally, unlike previous transfer-based attacks that treat each frame independently, overlooking the temporal dependencies between sequences, we integrate the temporal motion gradient in a Bayesian manner to disrupt the spatial-temporal coherence of S-HAR models. For exhaustive evaluation, we build the first comprehensive robust S-HAR evaluation benchmark RobustBenchHAR. RobustBenchHAR consists of 7 S-HAR models with diverse GCN structures and latest Transformer structures, 10 attack methods, 3 datasets and 2 defense methods. Extensive experiments demonstrate the superiority and generalizability of TASAR.
2 Related Work
@@
Skeleton-Based Human Action Recognition. Early S-HAR research employed convolutional neural networks (CNNs) (Ali et al., 2023) and recurrent neural networks (RNNs) (Du et al., 2015) to extract motion features in the spatial and temporal domains. However, skeleton data as a topological graph challenges feature representation with traditional methods. Recent advances with graph convolutional networks (GCNs) (Kipf & Welling, 2016) have improved performance by modeling skeletons as topological graphs, with nodes corresponding to joints and edges to bones (Yan et al., 2018). Subsequent improvements in graph designs and network architectures include two-stream adaptive GCN (2s-AGCN) (Shi et al., 2019a), directed acyclic GCN (DGNN) (Shi et al., 2019b), multi-scale GCN (MS-G3D) (Liu et al., 2020b), channel-wise topology refinement (CTR-GCN) (Chen et al., 2021) and auxiliary feature refinement (FR-HEAD) (Zhou et al., 2023). Alongside advancements in GCN-based models, recent studies have explored temporal Transformer structures for S-HARs (Do & Kim, 2024; Qiu et al., 2022; Guo et al., 2024), but their vulnerability remains unexplored. Recently, robust S-HAR against adversarial noise has gained attention, with works such as Diao et al. (2024a) exploring adversarial sample distributions and Tanaka et al. (2024) applying Fourier analysis. BEAT (Wang et al., 2023) employs a post-train Bayesian strategy to achieve full Bayesian treatment on clean data, adversarial distribution and classifier. Although post-train Bayesian strategy is suggested to be more robust (Wang et al., 2023), its application in S-HAR attacks has not been explored. To address this, we introduce a new post-train Dual Bayesian strategy to improve adversarial transferability.
@@
Adversarial Attacks on S-HAR. Adversarial attacks (Szegedy et al., 2013) have been applied to various data types, with increasing focus on S-HAR. CIASA (Liu et al., 2020a) proposes a constrained iterative attack via GAN (Goodfellow et al., 2014a) to regularize the adversarial skeletons, while SMART (Wang et al., 2021) uses a perception loss gradient. Tanaka et al. (2022) suggests only perturbing skeletal lengths. These methods are all white-box attacks, requiring full knowledge of the victim model. Different from existing white-box attacks leverage dynamics or physical constraints to preserve visual naturalness within white-box settings, we focus on disrupting spatial-temporal coherence to improve adversarial transferability. In contrast, BASAR (Diao et al., 2021; 2024a) proposes motion manifold searching to achieve the query-based black-box attack. FGDA-GS (Kang et al., 2023a) estimates gradient signs to further reduce query numbers. Compared to white-box and query-based attacks, transfer-based attacks (Liu et al., 2016) pose a more practical threat as real-world HAR scenarios typically cannot access white-box information or extensive querying. While existing white-box S-HAR attacks (Wang et al., 2021; Liu et al., 2020a) can be adapted for transfer-based scenarios, they suffer from low transferability and sensitivity to surrogate choices. Lu et al. (2023) proposes a no-box attack for S-HAR, but it also fails in transfer-based attacks. Various type of transfer-based attacks, including gradient-based (Dong et al., 2018; Ma et al., 2023; Ge et al., 2023), input transformation (Xie et al., 2019; Zhu et al., 2024; Wang et al., 2024), and ensemble-based methods (Xiong et al., 2022; Li et al., 2023; Tang et al., 2024), exhibit high transferability across various tasks but struggle in skeletal data (Lu et al., 2023). Therefore, there is an urgent need to develop a transferable attack for skeleton-based action recognition.
3 Methodology
3.1 Preliminaries
@@
We denote a clean motion $\mathbf{x}\in\mathcal{X}$ and its corresponding label $y\in\mathcal{Y}$. Given a surrogate action recognizer $f_{\theta}$ parametrized by $\theta$, $f_{\theta}$ is trained to map a motion $\mathbf{x}$ to a predictive distribution $p(y\mid\mathbf{x},\theta)$. The white-box attack aims to find adversarial examples $\tilde{\mathbf{x}}$ within the neighborhood $\mathcal{B}_{\epsilon}(\mathbf{x})=\{\tilde{\mathbf{x}}:\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon\}$ that misleads the target model $f_{\theta}$:
$\mathop{\arg\min}\limits_{\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon}p(y\mid\tilde{\mathbf{x}},\theta),$ (1)
@@
where $\epsilon$ is the perturbation budget. $\|\cdot\|_{p}$ is the $l_{p}$ norm distance. The procedure of transfer-based attack is firstly crafting the adversarial example $\tilde{\mathbf{x}}$ by attacking the surrogate model, then transferring $\tilde{\mathbf{x}}$ to attack the unseen target model. In Equation 1, since the transferable adversarial examples are optimized against one surrogate model, the adversarial transferability heavily relies on the surrogate model learning a classification boundary similar to that of the unknown target model. While possible for image classification, it proves unrealistic for S-HAR (Wang et al., 2023; Lu et al., 2023).
3.2 Motivation
@@
Existing S-HAR attacks have shown outstanding white-box attack performance but exhibit low transferability (Wang et al., 2023). Similarly, previous transfer-based attacks (Dong et al., 2018; Xiong et al., 2022), successful on image data, also show poor transferability when applied to skeletal motion (Lu et al., 2023). Naturally, two questions occur to us: (1) Why do existing adversarial attacks fail to exhibit transferability in skeletal data? (2) Do transferable adversarial examples truly exist in S-HAR?
@@
To answer these questions, we start by generating adversarial examples using various surrogate skeletal recognizers and then evaluate their adversarial transferability. Obviously, in Table 1, the transferability is highly sensitive to the chosen surrogates, e.g. CTR-GCN (Chen et al., 2021) as the surrogate exhibits higher transferability than ST-GCN (Yan et al., 2018). This observation motivates us to further investigate the differences between surrogate models. Previous research (Wu & Zhu, 2020; Qin et al., 2022) has proven that adversarial examples generated by surrogate models with a less smooth loss landscape are unlikely to transfer across models. Therefore, we investigate the smoothness of the loss landscape across different surrogate models. In Figure 2, we visualize the loss landscape of ST-GCN and CTR-GCN trained on the skeletal dataset NTU-60 (Shahroudy et al., 2016), and compare their smoothness to the ResNet-18 (He et al., 2016) trained on CIFAR-10 (Krizhevsky et al., 2009). More landscape visualizations can be found in Appendix C. By analyzing the loss surface smoothness, we have two findings: (1) The loss surface of models trained on skeletal data is much sharper than those trained on image data, leading to a relatively low transferability. This suggests that adversarial examples within a sharp local region are less likely to transfer across models in S-HAR, potentially explaining our first question. (2) CTR-GCN has a flatter loss landscape compared to ST-GCN, making it a more effective surrogate for higher transferability. Consequently, we argue that using a surrogate with a smoothed loss landscape will significantly enhance adversarial transferability in S-HAR.
@@
In this work, motivated by evidence that Bayesian neural networks (BNNs) exhibit low sharpness and good generalization (Blundell et al., 2015; Maddox et al., 2019), we aim to construct a Bayesian surrogate by sampling from the model posterior space to smoothen the rugged loss landscape. From a Bayesian perspective, Equation 1 can be reformulated by approximately minimizing the Bayesian posterior predictive distribution:
@@
$\underset{\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon}{\arg\min}\,p(y\mid\tilde{\mathbf{x}},\mathcal{D})=\underset{\|\tilde{\mathbf{x}}-\mathbf{x}\|_{p}\leq\epsilon}{\arg\min}\,\mathbb{E}_{\theta\sim p(\theta\mid\mathcal{D})}p\left(y\mid\tilde{\mathbf{x}},\theta\right),$ (2)
@@
Figure 2: Comparison of loss landscapes of trained models.The $x$ and $y$ axis represent two random direction vectors sampled from a Gaussian distribution, which are added to the model’s parameter space along these directions. These random direction vectors are used to assess the sensitivity of the model’s loss function. The $z$ axis represents the loss value. More details can be found in Li et al. (2018). BA means the Bayesian Attack proposed by Li et al. (2023). PB means the post-train Bayesian optimization, and P-DB means the improved post-train Dual Bayesian optimization. The loss landscape optimized by post-train Dual Bayesian is significantly smoother than those of vanilla post-train Bayesian and baseline methods. More visualizations can be found in Appendix C.
3.3 A Post-train Bayesian Perspective on Attack
Unfortunately, directly sampling from the posterior distribution of skeletal classifiers is not a straightforward task due to several factors. First, directly sampling the posterior is intractable for large-scale skeletal classifiers. Although approximate methods such as MCMC sampling (Welling & Teh, 2011) or variational inference (Blei et al., 2017) are possible, sampling is prohibitively slow and resource-intensive due to the high dimensionality of the sampling space, which typically involves at least several million parameters in skeletal classifiers. In addition, skeletal classifiers normally contain a large number of parameters and are pre-trained on large-scale datasets (Liu et al., 2019). Consequently, it is not practical for end-users to re-train the surrogate in a Bayesian manner, as the training process is time-consuming.